In 15 years, Maya will be twice as old as David is now. In 15 years, David will be as old as Maya will be in 10 years from now. how old are thet now?: In 15 years, Maya will be twice as old as David is now. In 15 years, David will be as old as Maya will be in 10 years from now. how old are thet now?
Answer by ankor@dixie-net.com(4508)

Let m = Maya's age now
Let d = David's age now
;
Write an equation for each statement;
"In 15 years, Maya will be twice as old as David is now."
m + 15 = 2d
:
"In 15 years, David will be as old as Maya will be in 10 years from now."
d + 15 = m + 10
d = m + 10 - 15
d = m - 5
:
how old are they now?
;
Substitute (m-5) for d in the 1st equation
m + 15 = 2(m-5)
m + 15 = 2m - 10
15 + 10 = 2m - m
m = 25 yrs is m's present age
:
d = m - 5
d = 25 - 5
d = 20 is d's present age
;
:
Check solutions in the statement:
"In 15 years, David will be as old as Maya will be in 10 years from now."
d + 15 = m + 10
20 + 15 = 25 + 10; confirms our solutions